OpenStudy (anonymous):

the change in an object's momentum is equal to A. the product of the mass of the object and the time ineterval B. the product of the force applied to the object and the time interval C. the time interval divided by the net external force D the net external force divided by the time interval

OpenStudy (anonymous):

@amistre64 @ajprincess @abb0t @agent0smith @sarahc @deezgurls @ganeshie8 @djcool31 @galacticwavesXX @jamie2012 @jamie2012 @Katie_405 @Luigi0210 @zkrup @Christos @BulletWithButterflyWings @Blaze @Brookebrookiee does anyone know this ?

OpenStudy (amistre64):

a change in **** tends to refer to the slope of a line; a derivative

OpenStudy (amistre64):

if i recall the setup: p = mv right?

OpenStudy (anonymous):

im pretty sure im just not sure what the answer is i thought it was answer B but im not for sure :P

ganeshie8 (ganeshie8):

you're saying Force x time = mv2 - mv1

OpenStudy (amistre64):

F = ma F = m \(\Large \frac{\Delta v}{t}\) Ft = m \( \Delta v\)

OpenStudy (anonymous):

im still confused o.O

OpenStudy (amistre64):

what is the confusion?

OpenStudy (anonymous):

well i dont really get the change in an object s momentum would equal to

OpenStudy (amistre64):

momentum changes when velocity changes change in momentum = m(v1-v2) = m \(\Delta v\)

OpenStudy (amistre64):

using Newtons second law, we can equate F with a change in momentum

OpenStudy (anonymous):

so i know its not B its not so would it be D ?

OpenStudy (amistre64):

.... force(time) = mass(change in velocity) force(time) = change in momentum

OpenStudy (amistre64):

just trying to show how we get to b, as opposed to just say that it is b :/

OpenStudy (anonymous):

okay thank you for the help :)

OpenStudy (amistre64):

good luck ;)